OpenStudy (anonymous):

I need help with this problem please! 5k^2+7k-5=0 solve using quadratic equation

OpenStudy (anonymous):

Jimmyrep help me please??

OpenStudy (anonymous):

u can solve this using quadratic formula k = -b +- sqrt(b^2- jim has done this for you

jimthompson5910 (jim_thompson5910):

\[\Large k = \frac{-b\pm\sqrt{b^2-4ac}}{2a} \] \[\Large k = \frac{-(7)\pm\sqrt{(7)^2-4(5)(-5)}}{2(5)} \] \[\Large k = \frac{-7\pm\sqrt{49-(-100)}}{10} \] \[\Large k = \frac{-7\pm\sqrt{149}}{10} \] \[\Large k = \frac{-7+\sqrt{149}}{10} \ \text{or} \ k = \frac{-7-\sqrt{149}}{10} \] \[\Large k = \frac{-7+\sqrt{149}}{10} \ \text{or} \ k = \frac{-7-\sqrt{149}}{10} \]